# KIX: A Knowledge and Interaction-Centric Metacognitive Framework for Task Generalization

### v2

@@ 3 Model
$\mathcal{L}_{p}^{\text{critic}}(\phi_{p})=\mathbb{E}_{\mathcal{D}_{\tau_{p}}}\left[\frac{1}{2}(V_{\mu_{\phi_{p}}}(s)-R_{p})^{2}\right]$
@@
$\mathcal{L}_{p}(\phi_{p})=\mathcal{L}_{p}^{\text{actor}}(\phi_{p})-\zeta_{p}^{h}\mathcal{H}_{\phi_{p}}+\zeta_{p}^{c}\mathcal{L}_{p}^{\text{critic}}(\phi_{p})$
@@
In summary, when viewing tasks as meta-policy guided interactions with objects, problems are naturally decomposed into interaction goals. Unlike the models that are tightly coupled with low-level representations through end-to-end policies, in the proposed framework, agent actions are aided by meta-level knowledge guided object interactions, providing a mechanism for generalist behaviors.
4 Results and Discussion
@@
Our goal is to examine whether meta-level knowledge-guided interactions benefit agents in task generalization. To assess the capabilities of the proposed agents, we need a challenging environment having ingredients for higher-order reasoning to solve tasks. MiniGrid world environments Chevalier-Boisvert et al. (2018) are suitable for this purpose; they are procedurally generated environments where the agent interacts with objects such as doors, keys, balls, and boxes.
ObstructedMaze-Full
ObstructedMaze-1Dlhb
@@
The obstructed maze environments, shown in Figure 3, are particularly challenging due to partial observability, stochasticity, and sparse rewards, with only a few claims for solving it. In the MiniGrid-ObstructedMaze-Full environment, the agent is initially placed in the middle room, with a 7x7 view. At the low level, the agent can move forward, left, right, pick, drop, and toggle, and its goal is to pick up the blue ball, which is randomly placed in one of the corner rooms. Some of the doors are locked and require a key of the same color to open. Keys are hidden inside boxes, but the agent cannot see inside the boxes without toggling. The default task is referred to as Task 0. Furthermore, we identify the following new task scenarios, which were presented to the agent during testing but not during training:
@@
Agents learning to interact using type space knowledge
@@
       /* collect exp for meta exp size K */
@@
We devised two variants: KIX-A and KIX-R. Both variants learn meta-policy and a collection of interaction policies. While the KIX-A variant combines both reaching objects and interacting with them, the KIX-R variant learns an additional interaction policy referred to as the ‘reachability’ policy, which is used to reach objects prior to interacting with them. All policies, including meta and interactions, are implemented as actor-critic PPO policies. Algorithm 1 provides an overview of the procedure. The agent’s partially observed view $X$ contains objects with identifiers and attributes like color and type. The function $\Phi$ takes $X$ and creates an instance space graph by using visible objects as nodes, including the agent, and determining relationships between nodes such as ‘visible’, ‘adjacent’, and ‘carrying’. The ‘adjacent’ relation is derived using distance between the two objects. When the agent chooses an object to interact with, an ‘activated’ relation is added between the agent and the object node. The function $\Psi$ then constructs a type space graph from the instance space graph, encoding node types as attribute vectors. Edges in a type space graph have the same meaning as those in an instance space graph and are encoded as one-hot encoded edge attribute vectors.
Figure 4: Agents learning to interact: The learning progress of the agents KIX-A (blue) and KIX-R (red) is shown as returns over steps, which illustrate both the meta-level and interactions such as ‘open’, ‘open with key’, ‘reveal’, ‘pickup’ (not shown for KIX-R), ‘drop’, and ‘reach’ in the ObstructedMaze-Full environment.
@@
The environment rewards the agents with $r=1-(c/c_{T})$ where $c$ is the number of steps taken and $c_{T}$ is the max possible steps. For interaction policies, the agent receives a reward of 1 when it has a successful interaction. At the meta-level, it is rewarded with 0.1 for a successful proposed interaction and $1+r$ for achieving the environment goal. The base agent, BASE, has a similar actor-critic architecture and receives environment rewards. The training hyper-parameters are given in Table 1. Learning rates were determined using grid search with values 1e-3, 1e-4, 1e-5, and 1e-6. All the agents are trained for $\sim 6.2\times 10^{8}$ steps, using 32 parallel workers.
As shown in Figure 4, as training progresses, agents learn to interact with objects and associate higher-level structured knowledge with these interactions using object types and knowledge graphs, as indicated by higher meta-level returns. This, in turn, leads to better interaction experiences and further improvements in interaction policies. Furthermore, the KIX-R variant first focuses on learning how to reach different objects, but once it learns to reach, it begins to learn interactions in fewer steps and with much higher returns compared to the KIX-A variant. It suggests that the agents learn higher-level relations and apply these learned relations and interactions with objects to solve problems.
Meta-level knowledge-guided interactions aid agents
@@
(b) Evaluation task returns
@@
The ObstructedMaze-1Dlhb environment is used to assess the transferability of interaction policies learned in one environment to another. In this environment, the agent’s goal is to pick up a blue ball hidden in a two-room maze. The rooms are separated by a door, which is obstructed by a ball and locked. The keys are hidden inside boxes. Agents employ previously acquired interaction policies from the base task in the ObstructedMaze-Full environment while only learning the meta-policies. During evaluation, agents use the learned meta-policy, but the interaction policies come from the ObstructedMaze-Full environment itself. Figure 6 shows the meta training returns and evaluation return profiles, similar to the main environment, for both agent variants. It is evident that agents have gained the ability to exploit interaction policies from the previous environment, and the second variant, which employs a separate reachability policy, appears to have benefited more.
5 Conclusion
In this work, we proposed a metacognitive framework that promotes generalist behaviors through three main contributions. First, generalist behaviors necessitate abstraction and composition; however, current models lack natural abstractions and compositions that can preserve both object abstractions and relational knowledge structures. In our model, type space knowledge graphs are used to represent and ground natural abstractions and compositions while also preserving relational knowledge between objects. Second, rather than directly acting in the environment, interactions with objects are critical in performing the tasks. To accomplish this, introduce an intermediary interaction level. The agent’s meta-level learns from type space knowledge graphs and makes recommendations on which objects to interact with and how to interact with them. In this sense, our model functions as an interaction goal-generating machine. Third, interactions employing interaction policies enable the necessary loose coupling required for generalization by acting as intermediaries. We demonstrated that metacognitive knowledge-guided interactions aid agents in generalization. Lastly, Our framework is modular and adaptable, allowing for the integration of knowledge structures and control variables at both the meta and interaction levels. This framework is promising, particularly for domains where higher-order knowledge and interaction compositions can be exploited, such as robotics, autonomous and cognitive systems.

### v3

@@ Introduction
$\mathcal{L}_{p}^{\text{critic}}(\phi_{p})=\mathbb{E}_{\mathcal{D}_{\tau_{p}}}\left[\frac{1}{2}(V_{\mu_{\phi_{p}}}(s)-R_{p})^{2}\right]$ (7)
@@
$\mathcal{L}_{p}(\phi_{p})=\mathcal{L}_{p}^{\text{actor}}(\phi_{p})-\zeta_{p}^{h}\mathcal{H}_{\phi_{p}}+\zeta_{p}^{c}\mathcal{L}_{p}^{\text{critic}}(\phi_{p})$ (8)
@@
In summary, when viewing tasks as meta-policy guided interactions with objects, problems are naturally decomposed into interaction goals. Unlike the models that are tightly coupled with low-level representations through end-to-end policies, in the proposed framework, agent actions are aided by the interpretation of meta-level knowledge to guide object interactions, providing a mechanism for generalist behaviors.
Results and discussion
@@
Our goal is to examine whether meta-level knowledge-guided interactions benefit agents in task generalization. To assess the capabilities of the proposed agents, we need a challenging environment having ingredients for higher-order reasoning to solve tasks. MiniGrid world navigation environments Chevalier-Boisvert et al. (2018) are suitable for this purpose; they are procedurally generated environments where the agent interacts with objects such as doors, keys, balls, and boxes. The obstructed maze environments, shown in Figure 3, are particularly challenging due to partial observability, stochasticity, and sparse rewards, with only a few claims for solving it. In the MiniGrid-ObstructedMaze-Full environment, the agent is initially placed in the middle room, with a 7x7 view. At the low level, the agent can move forward, left, right, pick, drop, and toggle, and its goal is to pick up the blue ball, which is randomly placed in one of the corner rooms. Some of the doors are locked and require a key of the same color to open. Keys are hidden inside boxes, but the agent cannot see inside the boxes without toggling.
(a) ObstructedMaze-Full
(b) ObstructedMaze-1Dlhb
@@
The default task is referred to as Task 0. Furthermore, we identify the following new task scenarios, which were presented to the agent during testing but not during training:
@@
Learning to interact using type space knowledge
We devised two variants: KIX-A and KIX-R. Both variants learn a meta-policy and a collection of interaction policies. While the KIX-A variant combines both reaching objects and interacting with them, the KIX-R variant learns an additional interaction policy referred to as the ‘reachability’ policy, which is used to reach objects prior to interacting with them. All policies, including meta and interactions, are implemented as actor-critic PPO policies.
@@
   /* collect meta experiences */
@@
Algorithm 1 provides an overview of the procedure. The agent’s partially observed view $X$ contains objects with identifiers and attributes like color and type. The function $\Phi$ takes $X$ and creates an instance space graph by using visible objects as nodes, including the agent, and determining relationships between nodes such as ‘visible’, ‘adjacent’, and ‘carrying’. The ‘adjacent’ relation is derived using distance between the two objects. When the agent chooses an object to interact with, an ‘activated’ relation is added between the agent and the object node. The function $\Psi$ then constructs a type space graph from the instance space graph, encoding node types as attribute vectors. Edges in a type space graph have the same meaning as those in an instance space graph and are encoded as one-hot encoded edge attribute vectors.
@@
The environment rewards the agents with $r=1-(c/c_{T})$ where $c$ is the number of steps taken and $c_{T}$ is the max possible steps. For interaction policies, the agent receives a reward of 1 when it has a successful interaction. At the meta-level, it is rewarded with 0.1 for a successful proposed interaction and $1+r$ for achieving the environment goal. The base agent, BASE, is an end to end policy learner with a similar actor-critic architecture and receives environment rewards. The training hyper-parameters are given in Table 1. Learning rates were determined using grid search with values 1e-3, 1e-4, 1e-5, and 1e-6. All the agents are trained for $\sim 6.2\times 10^{8}$ steps, using 32 parallel workers in a high performance computing cluster.
Figure 4: Agents learning to interact: The learning progress of the agents KIX-A (blue) and KIX-R (red) is shown as returns over steps, which illustrate learning both the meta-level and interactions such as ‘reach’ (only applicable for KIX-R), ‘open’, ‘open with key’, ‘reveal’, ‘pickup’, ‘drop’ (not shown) in the ObstructedMaze-Full environment.
As shown in Figure 4, as training progresses, agents learn to interact with objects and associate higher-level structured knowledge with these interactions using object types and type and instance space graphs, as indicated by higher meta-level returns. This, in turn, leads to better interaction experiences and further improvements in interaction policies. Furthermore, the KIX-R variant first focuses on learning how to reach different objects, but once it learns to reach, it begins to learn interactions in fewer steps and with much higher returns compared to the KIX-A variant. It suggests that the agents learn higher-order relations and apply these learned relations and interactions with objects to solve problems.
Meta-level guided interactions aid agents
@@
(b) Evaluation of task returns
@@
The ObstructedMaze-1Dlhb environment is used to assess the transferability of interaction policies learned in one environment to another. In this environment, the agent’s goal is to pick up a blue ball hidden in a two-room maze. The rooms are separated by a door that is obstructed by a ball and locked. The keys are hidden inside boxes. Agents employ previously acquired interaction policies from the base task in the ObstructedMaze-Full environment while only learning the meta-policies. During evaluation, agents use the learned meta-policy, but the interaction policies are from the ObstructedMaze-Full environment itself. Figure 6 shows the meta training returns and evaluation return profiles, similar to the main environment, for both agent variants. It is evident that agents have gained the ability to exploit interaction policies from the previous environment, and the second variant, which employs a separate reachability policy, appears to have benefited more.
Conclusion
In this work, we proposed a metacognitive framework that promotes generalist behaviors through three main contributions. First, generalist behaviors necessitate abstraction and composition; however, current models lack natural abstractions and compositions that can preserve both object abstractions and relational knowledge structures. In our model, type-space graphs are used to represent and ground natural abstractions and compositions while also preserving relational knowledge between objects. Second, rather than acting directly in the environment, interactions with objects are critical in performing tasks. To accomplish this, we introduced an intermediary interaction level. The agent’s meta-level interprets abstract relational knowledge, learns a meta-policy using type-space graphs, and makes recommendations on which objects to interact with and how to interact with them. In this sense, our model functions as an interpretation-based interaction goal-generating machine. Third, interactions employing interaction policies enable the necessary loose coupling required for generalization by acting as intermediaries. We demonstrated that metacognitive knowledge-guided interactions aid agents in generalization. Lastly, Our framework is modular and adaptable, enabling the integration of knowledge structures and control variables at both the meta and interaction levels. This framework is promising, particularly for domains where higher-order knowledge and interaction compositions can be exploited, such as robotics, autonomous and cognitive systems.